A Complete Bibliography of Publications in the Journal of Statistical Planning and Inference:

1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.01

Title word cross-reference

2^{3} [BS79a]. 2^{m} [Shi79, SG79]. 2l + 1 [Shi79]. 2 × 2 [Ber78b, Man77]. 3^{m} [Kuw79]. 4 [Kuw79]. m [Sin77]. n [KT79]. Φ_{p} [GK79b, Pes78]. q [GK79b]. t [Pel78, WO78].

-ary [KT79]. -associate [Sin77]. -ball [GK79b]. -Designs [Pel78]. -optimal [Pes78]. -optimum [GK79b]. -statistic [WO78].

1-106 [Ano78d]. 1-107 [Ano77b]. 1-90 [Ano79c]. 107-210 [Ano78e]. 109-216 [Ano77c]. 179-286 [Ano79d].

211-314 [Ano78f]. 217-312 [Ano77d]. 287-398 [Ano79e]. 2X2 [Ber78a].

91-178 [Ano79f]. above [Joh78]. alias [Shi79]. allow [SG79]. among [Joh78]. analogy [Man77]. Analysis [McC79, SMWH79, Sha77]. ANOVA [GS79]. application [Anb77, Fra78]. applications [RY79]. approximation [Anb77]. approximations [Kha78]. arcs [Bar79a]. arrays [DV78, Kuw79, Muk78, SS79b, SS79c, Wei78, Whi78]. ary [KT79]. aspects [GT78]. assay [Anb77]. associate [Sin77]. associated [Bha79, Pel78]. Asymptotic [PS77]. Author [Ano78a, Ano78b, Ano79a]. average [Kha78].

Balanced [Kuw79, BS79b, BO79, CM78, Lem79, Shi79, Tya79, VB79, Whi78]. ball [GK79b]. balls [GK77b]. based [McC79, Oga77]. Bayes [Var78]. be [BK77, Bar79b]. Behrens [WO78]. below [Joh78]. Berkson [Bar79b, Bas79, CdK79]. best [BK77]. between [GS79]. binomial [Var78]. bio [Anb77]. bio-assay [Anb77]. block [BS79b, CM78, KT77a, KT77b, KT79, Muk78, Tya79]. blocks [KT77b]. Board [Ano77a, Ano78c, Ano79b]. bound [KT77b]. Bounds [DV78, Var78].

can [Bar79b]. case [Kre78, Sen79]. censored [McC79]. censoring [Joh78, Joh78]. certain [KT77a]. change [Wei78]. change-over [Wei78]. Characterization [KT77a]. Characterizations [ENG79]. class [BS79b, Tya79]. classification [Ban79, McC79]. collineation [HW79]. Comment [CdK79]. common [KT77b, Man77]. Comparison [GK77a, GK77b]. comparisons [Joh78]. competition [VB79]. complete [GK79a]. Completeness [Joh78]. compound [Sin78, Var78]. computer [Whi78]. conditional [Bar79b]. Congruence [DJ77]. conics [Pel78]. consecutive [YUT^{+}77]. Construction [AHL77, Che79, BM78, Jun79, Sin77]. contain [Ber78b, Ber78a]. contingency [Man77]. contradiction [Bar79b]. convergence [Sin78]. correlated [KKS77]. covariance [PS77]. covariances [Kha79]. criterion [BK77]. cubes [GK77a]. cubic [GK79b]. cusum [Kha78].

decomposition [RY79]. deficiency [SS79b]. definition [RY79]. degenerate [Pel78]. derivation [GS79]. design [CM78, GK77a]. Designs [Pel78, VB79, AH79, AHL77, BM78, Bos77, BS79b, Bus79, BO79, DJ77, GK79a, GK77b, GK79b, HS79, JT77, Jun79, KT77a, KT79, Kuw79, Muk78, PRW78, Pes78, Sha77, Shi79, Sin77, Tya79, Wei78]. DFR [ENG79]. dimensional [Dav78]. disclosure [Fra78]. discrete [ENG79]. discriminant [CS79, SC78]. Discussion [Bas79]. dispraise [Bar79b, Bas79, Ber78b, CdK79, Kem79]. distribution [ENG79, KKS77, McC79, Oga77, PS77, SC78]. distributional [Bha79]. distributions [ENG79, SS79a]. divisible [BM78, Bos77, Bus79, JT77, Jun79]. Do [Ber78b, Ber78a]. dual [Bos77, CM78].

Edgeworth [SC78]. Editor [Sri79]. Editorial [Ano77a, Ano78c, Ano79b]. effect [PRW78, SG79]. effects [GS79]. efficient [Bar79b]. elimination [Sha77]. Embedding [SS79b, SS79c]. empirical [Var78]. encountered [Dav78]. equality [OB79, PS77]. equations [GT78]. equireplicated [KT79]. estimate [Kha79]. estimating [GT78, Lem79]. Estimation [Bha79, Dav78, Oga77, RB78, Sin78, SG79]. estimator [BK77, Var78]. estimators [Kre78]. exact [Bas79, Ber78b, CdK79, DHS79, Kem79]. experiments [BS79a, McC79]. exponential [Sin78]. extended [Sen79]. extension [BK77]. Extrapolation [GK79b].

factorial [BS79a, Kuw79, Shi79]. factorials [SG79]. Families [Bus79, Sin78]. February [Ano77b]. finite [OB79, RB78]. Fisher [WO78]. fixed [GS79]. fixed-effects [GS79]. fluctuation [SS79a]. fold [PRW78]. fold-over [PRW78]. forms [KKS77]. fractional [Kuw79, Shi79]. function [CS79, Kha79, SC78].

Galois [Pel78]. Gaussian [Ban79]. General [KT77b, RY79, Joh78]. generalized [RB78]. generation [Whi78]. geometric [ENG79]. given [KT77b]. graphs [Fra77]. greater [SS79b]. group [Bos77, Bus79, JT77, Jun79]. groups [HW79].

Hadamard [Bus79]. Haenszel [Man77]. half [Lem79]. half-sample [Lem79]. having [KT77b, YUT^{+}77]. hazards [GKAH77]. heterogeneity [Sha77]. Hotelling [PS77]. housing [SMWH79]. hyperrectangles [Che79].

IFR [ENG79]. II [HW79, Joh78]. Incomplete [Muk78, BS79b, KT77a]. index [Ano78a, Ano78b, Ano79a]. industry [Dav78]. Inequality [KT79]. information [Ber78b, Ber78a, Dal77, Fra78]. intermediate [GS79]. irreducible [HW79].

J. [Bar79b, CdK79]. Johnson [CS79]. joint [KKS77, SS79a]. Joseph [Bas79].

Klingenberg [DJ77]. Knut [AHL77].

Latin [Dra77]. layouts [GK79a]. least [BK77]. length [Kha78]. limits [Man77]. linear [BK77, Bha79, CS79, Kha79, Kre78, SC78]. location [Oga77].

Main [SG79, PRW78]. Manova [Kha79]. Mantel [Man77]. marginal [Ber78b, Ber78a]. matrices [PS77, Shi79]. Maximal [Dra77]. mean [PS77, RB78]. method [Anb78, BM78]. methods [Anb77, Man77]. mineral [Dav78]. minimal [YUT^{+}77]. Minimum [Kha79]. Misclassification [BS79a]. model [GS79, GKAH77, Kha79]. models [GS79, RB78].

Nets [BO79]. Newton [Anb78]. non [Dav78, Pel78]. nonnormality [CS79, SC78]. nonparametric [Sen79]. norm [Shi79]. normal [Oga77]. note [KKS77, SS79c, Sri79]. number [KT77b, YUT^{+}77].

observation [VB79]. odd [Pel78]. odds [Man77]. one [SG79]. Optimal [Oga77, RB78, Pes78]. optimality [AH79]. Optimum [GK79a, GK79b]. order [ENG79, Pel78, Pes78]. orthogonal [Muk78, SS79b, SS79c].

Pages [Ano77b, Ano77c, Ano77d, Ano78d, Ano78e, Ano78f, Ano79c, Ano79f, Ano79d, Ano79e]. paper [Bas79, CdK79]. parameter [McC79, Var78]. parameters [Oga77]. partial [Bas78, DJ77, Dra77]. partially [BS79b, BO79]. Partition [YUT^{+}77]. PBIB [Sin77]. permutation [DV78, RB78]. Perspectivities [HW79]. plan [SG79]. plane [Pel78]. planes [HW79]. plant [VB79]. point [BM78]. points [Bar79a]. population [Ban79, RB78]. populations [OB79]. post [DHS79]. post-stratification [DHS79]. Power [GS79]. preliminary [WO78]. Privacy [Dal77]. probabilities [BS79a]. problem [Anb77, Bha79, Fra78, Sen79, WO78]. problems [Dav78, RY79]. procedure [Man77]. procedures [Kha78, WO78]. process [Ban79]. processes [Dav78]. projective [HW79]. projectors [RY79]. properties [CM78]. property [Bos77, YUT^{+}77]. proportions [Ber78b, Ber78a].

quadratic [GK77a, GK77b, KKS77, Kha79]. quality [SMWH79]. quantiles [Oga77]. query [YUT^{+}77]. question [GKAH77].

random [GS79, RB78]. random-effects [GS79]. randomize [Kem77]. randomized [GKAH77]. rank [BK77]. Raphson [Anb78]. ratio [Man77]. ratios [OB79]. reactions [Kem79]. regions [Pes78]. regression [GK77a, GK77b, GK79b]. relations [DJ77]. relevant [Ber78b, Ber78a]. reliability [SMWH79]. remarks [HS79]. replication [Kre78]. resolution [Shi79]. respecting [Ber78b, Ber78a]. Respondent [GKAH77]. response [AH79, GKAH77]. retrieval [YUT^{+}77]. review [Bas78]. risks [Var78]. Robbins [Var78]. Robustness [CS79, SC78, AH79, PS77]. rotatable [GK77b]. run [Kha78].

sample [DHS79, Lem79, Oga77, Sen79]. samples [Joh78, McC79]. Sampling [Dav78, Fra77, SS77]. scale [Oga77, SMWH79, Sin78]. scale-exponential [Sin78]. search [SG79]. second [Pes78]. selected [Oga77]. Sequence [Sin78]. Sequence-compound [Sin78]. sequences [Joh78]. Serial [Wei78]. series [SC78]. set [YUT^{+}77]. sets [Dra77]. several [Man77]. simultaneous [Oga77]. single [McC79]. sizes [KT79]. small [DHS79]. solutions [DJ77]. Some [GT78, HS79, JT77, CM78, Kre78, RY79, SS79a]. spacings [Oga77]. spatial [Dav78]. speed [Sin78]. squares [BK77, Dra77]. stability [Kre78]. stationary [Dav78]. statistic [WO78]. statistical [Dal77, Fra78, Lem79, RY79]. statistics [ENG79]. stochastic [Anb77, Anb78, Dav78]. stratification [DHS79]. strength [Kuw79, SS79b, SS79c]. structures [DJ77]. study [PS77]. subdesigns [HS79]. subsets [YUT^{+}77]. sufficiency [Bas78]. surface [AH79]. Survey [Fra77]. Symmetric [Bos77, HS79, Pes78]. system [CS79]. systematic [SS77]. systems [Dal77].

table [Ber78b, Ber78a]. tables [Man77]. technique [Lem79]. test [Bas79, Ber78b, CdK79, Kem79, PS77]. Testing [OB79, WO78]. Tests [Man77, Bar79b]. theory [DHS79, Fra78, GT78, SS79a]. three [Dav78]. totally [CM78]. totals [Ber78b, Ber78a]. trace [PS77]. transformation [Bha79]. transformations [Dal77]. transversals [Dra77]. treatments [KT77b]. truncated [Var78]. Two [Ban79, GK79a, McC79, Oga77, OB79, PS77, Sen79, Sha77, SS79b, SS79c]. two-parameter [McC79]. two-sample [Sen79]. two-way [GK79a, Sha77].

unbiased [BK77, Kha79]. unequal [KT79, PS77]. unique [Lem79]. unknown [SG79]. unrelated [GKAH77]. upper [KT77b]. use [Lem79, WO78]. using [ENG79].

variance [Kha79, Kre78, Tya79]. variances [Kha79, Lem79]. vectors [PS77]. versions [Var78]. Vik [AHL77]. Volume [Ano78b, Ano79a, Ano78a].

Wald [Kha78]. way [GK79a, Sha77]. Weibull [McC79]. weighted [Bar79a]. weights [Lem79]. which [GS79, SG79].

Youden [Che79].

References

Andrews:1979:ROR

[AH79] David F. Andrews and Agnes M. Herzberg. The robustness and optimality of response surface designs. Journal of Statistical Planning and Inference, 3(3):249–257, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

          sciencedirect.com/science/article/pii/0378375879900168.

Atkin:1977:CKV

[AHL77] A. O. L. Atkin, L. Hay, and R. G. Larson. Construction of Knut Vik designs. Journal of Statistical Planning and Inference, 1(3): 289–297, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print),

          1873-1171 (electronic). URL http://www.sciencedirect.com/
          science/article/pii/0378375877900131.

Anbar:1977:ASA

[Anb77] Dan Anbar. The application of stochastic approximation methods to the bio-assay problem. Journal of Statistical Planning and Inference, 1(2):191–206, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375877900234.

Anbar:1978:SNR

[Anb78] Dan Anbar. A stochastic Newton–Raphson method. Journal of Statistical Planning and Inference, 2(2):153–163, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375878900046.

Anonymous:1977:EB

[Ano77a] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 1(1):ii, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375877900015.

Anonymous:1977:PF

[Ano77b] Anonymous. Pages 1-107 (February 1977). Journal of Statistical Planning and Inference, 1(1):??, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1977:Pa

[Ano77c] Anonymous. Pages 109-216 (1977). Journal of Statistical Planning and Inference, 1(2):??, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1977:Pb

[Ano77d] Anonymous. Pages 217-312 (1977). Journal of Statistical Planning and Inference, 1(3):??, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1978:AIVa

[Ano78a] Anonymous. Author index, volume 1 (1977). Journal of Statistical Planning and Inference, 2(1):105–106, ???? 1978. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375878900277.

Anonymous:1978:AIVb

[Ano78b] Anonymous. Author index: Volume 2 (1978). Journal of Statistical Planning and Inference, 2(3):313–314, ???? 1978. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375878900150.

Anonymous:1978:EB

[Ano78c] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 2(1):ii, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375878900162.

Anonymous:1978:Pa

[Ano78d] Anonymous. Pages 1-106 (1978). Journal of Statistical Planning and Inference, 2(1):??, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1978:Pb

[Ano78e] Anonymous. Pages 107-210 (1978). Journal of Statistical Planning and Inference, 2(2):??, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1978:Pc

[Ano78f] Anonymous. Pages 211-314 (1978). Journal of Statistical Planning and Inference, 2(3):??, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1979:AIV

[Ano79a] Anonymous. Author index: Volume 3 (1979). Journal of Statistical Planning and Inference, 3(4):395–397, ???? 1979. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375879900363.

Anonymous:1979:EB

[Ano79b] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 3(1):ii, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375879900375.

Anonymous:1979:Pa

[Ano79c] Anonymous. Pages 1-90 (1979). Journal of Statistical Planning and Inference, 3(1):??, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1979:Pc

[Ano79d] Anonymous. Pages 179-286 (1979). Journal of Statistical Planning and Inference, 3(3):??, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1979:Pd

[Ano79e] Anonymous. Pages 287-398 (1979). Journal of Statistical Planning and Inference, 3(4):??, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1979:Pb

[Ano79f] Anonymous. Pages 91-178 (1979). Journal of Statistical Planning and Inference, 3(2):??, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Bandyopadhyay:1979:TPC

[Ban79] Shibdas Bandyopadhyay. Two population classification in Gaussian process. Journal of Statistical Planning and Inference, 3(3): 225–233, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900144.

Barnabei:1979:AWP

[Bar79a] Marilena Barnabei. On arcs with weighted points. Journal of Statistical Planning and Inference, 3(3):279–286, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837587990020X.

Barnard:1979:CJB

[Bar79b] G. A. Barnard. In contradiction to J. Berkson’s dispraise: conditional tests can be more efficient. Journal of Statistical Planning and Inference, 3(3):181–187, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375879900090.

Basu:1978:PSR

[Bas78] D. Basu. On partial sufficiency: a review. Journal of Statistical Planning and Inference, 2(1):1–13, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375878900174.

Basu:1979:DJB

[Bas79] D. Basu. Discussion of Joseph Berkson’s paper “In dispraise of the exact test”. Journal of Statistical Planning and Inference, 3(3): 189–192, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900107. See [Ber78b].

Berkson:1978:DMT

[Ber78a] Joseph Berkson. Do the marginal totals of the 2x2 table contain relevant information respecting the table proportions? Journal of Statistical Planning and Inference, 2(1):43–44, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375878900204.

Berkson:1978:DET

[Ber78b] Joseph Berkson. In dispraise of the exact test: Do the marginal totals of the 2 × 2 table contain relevant information respecting the table proportions? Journal of Statistical Planning and Inference, 2(1):27–42, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375878900198. See comments

[Bas79, CdK79, Kem79].

Bhargava:1979:ELT

[Bha79] Anil K. Bhargava. Estimation of a linear transformation and an associated distributional problem. Journal of Statistical Planning and Inference, 3(1):19–26, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375879900399.

Baksalary:1977:ERC

[BK77] J. K. Baksalary and R. Kala. An extension of a rank criterion for the least squares estimator to be the best linear unbiased estimator. Journal of Statistical Planning and Inference, 1(3):309– 312, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375877900155.

Beker:1978:CMP

[BM78] Henry J. Beker and Christopher J. Mitchell. A construction method for point divisible designs. Journal of Statistical Planning and Inference, 2(3):293–306, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375878900137.

Bush:1979:NPB

[BO79] K. A. Bush and T. G. Ostrom. Nets and partially balanced designs. Journal of Statistical Planning and Inference, 3(4):381– 385, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837587990034X.

Bose:1977:SGD

[Bos77] R. C. Bose. Symmetric group divisible designs with the dual property. Journal of Statistical Planning and Inference, 1(1):87– 101, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375877900088.

Bohrer:1979:MPF

[BS79a] Robert Bohrer and Judith Sheft. Misclassification probabilities in 2^{3} factorial experiments. Journal of Statistical Planning and Inference, 3(1):79–84, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375879900430.

Bose:1979:CPB

[BS79b] R. C. Bose and M. S. Shrikhande. On a class of partially balanced incomplete block designs. Journal of Statistical Planning and Inference, 3(2):91–99, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375879900016.

Bush:1979:FHG

[Bus79] K. A. Bush. Families of Hadamard group divisible designs. Journal of Statistical Planning and Inference, 3(4):387–394, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900351.

Corsten:1979:CJB

[CdK79] L. C. A. Corsten and J. P. M. de Kroon. Comment on J. Berkson’s paper “In dispraise of the exact test”. Journal of Statistical Planning and Inference, 3(3):193–197, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900119. See [Ber78b].

Cheng:1979:CYH

[Che79] Ching-Shui Cheng. Construction of Youden hyperrectangles. Journal of Statistical Planning and Inference, 3(2):109–118, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837587990003X.

Ceranka:1978:SPD

[CM78] Bronislaw Ceranka and Stanislaw Mejza. On some properties of the dual of a totally balanced block design. Journal of Statistical Planning and Inference, 2(1):93–94, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375878900253.

Chinganda:1979:RLD

[CS79] Enock F. Chinganda and Kocherlakota Subrahmaniam. Robustness of the linear discriminant function to nonnormality: Johnson’s system. Journal of Statistical Planning and Inference, 3(1): 69–77, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900429.

Dalenius:1977:PTS

[Dal77] Tore Dalenius. Privacy transformations for statistical information systems. Journal of Statistical Planning and Inference, 1(1):73– 86, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375877900076.

David:1978:SEP

[Dav78] Michel David. Sampling and estimation problems for three dimensional spatial stationary and non stationary stochastic processes as encountered in the mineral industry. Journal of Statistical Planning and Inference, 2(3):211–244, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375878900095.

Doss:1979:ESS

[DHS79] D. C. Doss, H. O. Hartley, and G. R. Somayajulu. An exact small sample theory for post-stratification. Journal of Statistical Planning and Inference, 3(3):235–247, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375879900156.

Drake:1977:KSP

[DJ77] David A. Drake and Dieter Jungnickel. Klingenberg structures and partial designs i: Congruence relations and solutions. Journal of Statistical Planning and Inference, 1(3):265– 287, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837587790012X.

Drake:1977:MSL

[Dra77] David A. Drake. Maximal sets of Latin squares and partial transversals. Journal of Statistical Planning and Inference, 1(2): 143–149, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375877900192.

Deza:1978:BPA

[DV78] M. Deza and S. A. Vanstone. Bounds for permutation arrays. Journal of Statistical Planning and Inference, 2(2):197– 209, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375878900083.

El-Neweihi:1979:CGD

[ENG79] Emad El-Neweihi and Z. Govindarajulu. Characterizations of geometric distribution and discrete IFR (DFR) distributions using order statistics. Journal of Statistical Planning and Inference, 3 (1):85–90, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900442.

Frank:1977:SSG

[Fra77] Ove Frank. Survey sampling in graphs. Journal of Statistical Planning and Inference, 1(3):235–264, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375877900118.

Frank:1978:AIT

[Fra78] Ove Frank. An application of information theory to the problem of statistical disclosure. Journal of Statistical Planning and Inference, 2(2):143–152, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375878900034.

Galil:1977:CDQ

[GK77a] Z. Galil and J. Kiefer. Comparison of design for quadratic regression on cubes. Journal of Statistical Planning and Inference, 1(2):121–132, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375877900179.

Galil:1977:CRD

[GK77b] Z. Galil and J. Kiefer. Comparison of rotatable designs for regression on balls, I (quadratic). Journal of Statistical Planning and Inference, 1(1):27–40, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375877900039.

Gaffke:1979:ODC

[GK79a] N. Gaffke and O. Krafft. Optimum designs in complete two-way layouts. Journal of Statistical Planning and Inference, 3(2):119– 126, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900041.

Galil:1979:EDO

[GK79b] Z. Galil and J. Kiefer. Extrapolation designs and Φ_{p}-optimum designs for cubic regression on the q-ball. Journal of Statistical Planning and Inference, 3(1):27–38, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375879900405.

Greenberg:1977:RHU

[GKAH77] B. G. Greenberg, R. R. Kuebler, J. R. Abernathy, and D. G. Horvitz. Respondent hazards in the unrelated question randomized response model. Journal of Statistical Planning and Inference, 1(1):53–60, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375877900052.

Genizi:1979:PDA

[GS79] A. Genizi and M. Soller. Power derivation in an ANOVA model which is intermediate between the “fixed-effects” and the “randomeffects” models. Journal of Statistical Planning and Inference, 3(2): 127–134, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900053.

Godambe:1978:SAT

[GT78] V. P. Godambe and M. E. Thompson. Some aspects of the theory of estimating equations. Journal of Statistical Planning and Inference, 2(1):95–104, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375878900265.

Haemers:1979:SRS

[HS79] Willem Haemers and Mohan Shrikhande. Some remarks on subdesigns of symmetric designs. Journal of Statistical Planning and Inference, 3(4):361–366, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375879900326.

Hering:1979:PIC

[HW79] Christoph Hering and Michael Walker. Perspectivities in irreducible collineation groups of projective planes, II. Journal of Statistical Planning and Inference, 3(2):151–177, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900077.

Johnson:1978:CCA

[Joh78] N. L. Johnson. Completeness comparisons among sequences of samples II. Censoring from above or below, and general censoring. Journal of Statistical Planning and Inference, 2(2):107– 123, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375878900010.

John:1977:SNG

[JT77] J. A. John and G. Turner. Some new group divisible designs. Journal of Statistical Planning and Inference, 1(1):103– 107, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837587790009X.

Jungnickel:1979:CGD

[Jun79] Dieter Jungnickel. A construction of group divisible designs. Journal of Statistical Planning and Inference, 3(3):273–278, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900193.

Kempthorne:1977:WR

[Kem77] Oscar Kempthorne. Why randomize? Journal of Statistical Planning and Inference, 1(1):1–25, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375877900027.

Kempthorne:1979:DET

[Kem79] Oscar Kempthorne. In dispraise of the exact test: reactions. Journal of Statistical Planning and Inference, 3(3):199–213, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900120. See [Ber78b].

Khan:1978:WAA

[Kha78] Rasul A. Khan. Wald’s approximations to the average run length in cusum procedures. Journal of Statistical Planning and Inference, 2(1):63–77, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/037837587890023X.

Khatri:1979:MVQ

[Kha79] C. G. Khatri. Minimum variance quadratic unbiased estimate of a linear function of variances and covariances under manova model. Journal of Statistical Planning and Inference, 3(4):299– 303, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900259.

Khatri:1977:NJD

[KKS77] C. G. Khatri, P. R. Krishnaiah, and P. K. Sen. A note on the joint distribution of correlated quadratic forms. Journal of Statistical Planning and Inference, 1(3):299–307, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375877900143.

Krewski:1978:SSR

[Kre78] Daniel Krewski. On the stability of some replication variance estimators in the linear case. Journal of Statistical Planning and Inference, 2(1):45–51, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375878900216.

Kageyama:1977:CCI

[KT77a] Sanpei Kageyama and Takumi Tsuji. Characterization of certain incomplete block designs. Journal of Statistical Planning and Inference, 1(2):151–161, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375877900209.

Kageyama:1977:GUB

[KT77b] Sanpei Kageyama and Takumi Tsuji. General upper bound for the number of blocks having a given number of treatments common with a given block. Journal of Statistical Planning and Inference, 1(3):217–234, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375877900106.

Kageyama:1979:IEA

[KT79] Sanpei Kageyama and Takumi Tsuji. Inequality for equireplicated n-ary block designs with unequal block sizes. Journal of Statistical Planning and Inference, 3(2):101–107, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900028.

Kuwada:1979:BAS

[Kuw79] Masahide Kuwada. Balanced arrays of strength 4 and balanced fractional 3^{m} factorial designs. Journal of Statistical Planning and Inference, 3(4):347–360, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375879900314.

Lemeshow:1979:UUS

[Lem79] Stanley Lemeshow. The use of unique statistical weights for estimating variances with the balanced half-sample technique. Journal of Statistical Planning and Inference, 3(4):315–323, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900272.

Mantel:1977:TLC

[Man77] Nathan Mantel. Tests and limits for the common odds ratio of several 2 × 2 contingency tables: methods in analogy with the Mantel– Haenszel procedure. Journal of Statistical Planning and Inference, 1(2):179–189, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375877900222.

McCool:1979:ASC

[McC79] J. I. McCool. Analysis of single classification experiments based on censored samples from the two-parameter Weibull distribution. Journal of Statistical Planning and Inference, 3(1):39– 68, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900417.

Mukhopadhyay:1978:IBD

[Muk78] A. C. Mukhopadhyay. Incomplete block designs through orthogonal and incomplete orthogonal arrays. Journal of Statistical Planning and Inference, 2(2):189–195, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375878900071.

Otis:1979:TER

[OB79] David Otis and David C. Bowden. Testing equality of ratios from two finite populations. Journal of Statistical Planning and Inference, 3(2):135–150, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375879900065.

Ogawa:1977:OSS

[Oga77] Junjiro Ogawa. Optimal spacings for the simultaneous estimation of the location and scale parameters of a normal distribution based on selected two sample quantiles. Journal of Statistical Planning and Inference, 1(1):61–72, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375877900064.

Pellegrino:1978:DAN

[Pel78] Giuseppe Pellegrino. t-designs associated with non degenerate conics in a Galois plane of odd order. Journal of Statistical Planning and Inference, 2(3):307–312, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375878900149.

Pesotchinsky:1978:OSO

[Pes78] L. Pesotchinsky. Φ_{p}-optimal second order designs for symmetric regions. Journal of Statistical Planning and Inference, 2(2): 173–188, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837587890006X.

Pesotan:1978:MEF

[PRW78] H. Pesotan, B. L. Raktoe, and R. G. Worthley. On main effect fold-over designs. Journal of Statistical Planning and Inference, 2(3):277–291, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375878900125.

Pillai:1977:ADH

[PS77] K. C. S. Pillai and Nashat B. Saweris. Asymptotic distribution of Hotelling’s trace for two unequal covariance matrices and robustness study of test of equality of mean vectors. Journal of Statistical Planning and Inference, 1(2):109–120, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375877900167.

Rao:1978:OEF

[RB78] J. N. K. Rao and D. R. Bellhouse. Optimal estimation of a finite population mean under generalized random permutation models. Journal of Statistical Planning and Inference, 2(2):125– 141, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375878900022.

Rao:1979:GDD

[RY79] C. Radhakrishna Rao and Haruo Yanai. General definition and decomposition of projectors and some applications to statistical problems. Journal of Statistical Planning and Inference, 3(1): 1–17, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900387.

Subrahmaniam:1978:RLD

[SC78] Kocherlakota Subrahmaniam and Enock F. Chingánda. Robustness of the linear discriminant function to nonnormality: Edgeworth series distribution. Journal of Statistical Planning and Inference, 2(1):79–91, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375878900241.

Sen:1979:ETS

[Sen79] Pranab Kumar Sen. The extended two-sample problem: nonparametric case. Journal of Statistical Planning and Inference, 3(4): 287–298, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900247.

Srivastava:1979:MEP

[SG79] J. N. Srivastava and B. C. Gupta. Main effect plan for 2^{m} factorials which allow search and estimation of one unknown effect. Journal of Statistical Planning and Inference, 3(3):259– 265, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837587990017X.

Shah:1977:ADT

[Sha77] K. R. Shah. Analysis of designs with two-way elimination of heterogeneity. Journal of Statistical Planning and Inference, 1(2): 207–216, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375877900246.

Shirakura:1979:NAM

[Shi79] Teruhiro Shirakura. On the norm of alias matrices in balanced fractional 2^{m} factorial designs of resolution 2l + 1. Journal of Statistical Planning and Inference, 3(4):337–345, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900302.

Sinha:1977:CAP

[Sin77] Kishore Sinha. On the construction of m-associate PBIB designs I. Journal of Statistical Planning and Inference, 1(2):133– 142, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375877900180.

Singh:1978:SCE

[Sin78] R. S. Singh. Sequence-compound estimation in scale-exponential families and speed of convergence. Journal of Statistical Planning and Inference, 2(1):53–62, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375878900228.

Schucany:1979:ARN

[SMWH79] William R. Schucany, Edward R. Mansfield, Wayne A. Woodward, and James L. Hess. Analysis of the reliability of a new scale for housing quality. Journal of Statistical Planning and Inference, 3 (4):305–313, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375879900260.

Srivastava:1979:EN

[Sri79] J. N. Srivastava. Editor’s note. Journal of Statistical Planning and Inference, 3(3):179, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375879900089.

Singh:1977:NSS

[SS77] D. Singh and Padam Singh. New systematic sampling. Journal of Statistical Planning and Inference, 1(2):163–177, ???? 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375877900210.

Saran:1979:SJD

[SS79a] Jagdish Saran and Kanwar Sen. On some joint distributions in fluctuation theory. Journal of Statistical Planning and Inference, 3(3):215–223, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375879900132.

Shrikhande:1979:EOA

[SS79b] S. S. Shrikhande and N. M. Singhi. Embedding of orthogonal arrays of strength two and deficiency greater than two. Journal of Statistical Planning and Inference, 3(4):367–379, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900338.

Shrikhande:1979:NEO

[SS79c] S. S. Shrikhande and N. M. Singhi. A note on embedding of orthogonal arrays of strength two. Journal of Statistical Planning and Inference, 3(3):267–271, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375879900181.

Tyagi:1979:CVB

[Tya79] B. N. Tyagi. On a class of variance balanced block designs. Journal of Statistical Planning and Inference, 3(4):333–336, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375879900296.

Vardeman:1978:BEB

[Var78] Stephen B. Vardeman. Bounds on the empirical Bayes and compound risks of truncated versions of Robbins’s estimator of a binomial parameter. Journal of Statistical Planning and Inference, 2(3):245–252, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375878900101.

Veevers:1979:DBO

[VB79] A. Veevers and T. B. Boffey. Designs for balanced observation of plant competition. Journal of Statistical Planning and Inference, 3(4):325–331, ???? 1979. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375879900284.

Weidman:1978:SAC

[Wei78] Lynn Weidman. Serial arrays and change-over designs. Journal of Statistical Planning and Inference, 2(2):165–171, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375878900058.

White:1978:CGB

[Whi78] David White. On computer generation of balanced arrays. Journal of Statistical Planning and Inference, 2(3):253–275, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375878900113.

Wehrhahn:1978:USP

[WO78] K. Wehrhahn and J. Ogawa. On the use of the t-statistic in preliminary testing procedures for the Behrens–Fisher problem. Journal of Statistical Planning and Inference, 2(1):15– 25, ???? 1978. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375878900186.

Yamamoto:1977:PQS

[YUT^{+}77] Sumiyasu Yamamoto, Kazuhiko Ushio, Shinsei Tazawa, Hideto Ikeda, Fumikazu Tamari, and Noboru Hamada. Partition of a query set into minimal number of subsets having consecutive retrieval property. Journal of Statistical Planning and Inference, 1(1):41–51, February 1977. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0378375877900040.